A net force of –8750N is used to stop of 1250.kg car travelling 25m/s. What braking distance is needed to bring the car to a halt?


1
Expert's answer
2021-03-29T09:58:43-0400

Let's first find the deceleration of the car:


"a=\dfrac{F}{m}=\dfrac{–8750\ N}{1250\ kg}=-7\ \dfrac{m}{s^2}."

Finally, we can find the braking distance needed to bring the car to a halt:


"v^2=v_0^2+2as,""0=v_0^2+2as,""s=\dfrac{-v_0^2}{2a}=\dfrac{-(25\ \dfrac{m}{s})^2}{2\cdot(-7\ \dfrac{m}{s^2})}=44.64\ m."